Suppose that the economy is at full-employment equilibrium.
(a) Explain what the Keynesian multiplier effect is in macroeconomics. Using the Keynesian
cross, explain mathematically why the Keynesian multiplier effect is greater than 1.
(b) In the Keynesian cross, suppose that the consumption function is given by:
C = 100 + 0.6 x (Y – T).
Also suppose that the planned investment (I) is 50 and the government expenditure and
taxes are both 200. Assume that there is no international trade (both exports and imports
are zero).
(i) Calculate the equilibrium level of GDP (income).
(ii) Calculate the equilibrium level of GDP (income) when the government decreases its
expenditure from 200 to 150.
(iii) Suppose that the government attempts to achieve an output (GDP) of 1,000. What level
of government expenditure is required?
1
Expert's answer
2017-04-14T09:11:05-0400
(a) Keynesian multiplier effect is one of the chief components of Keynesian economic models. According to Keynes theory of fiscal stimulus, an injection of government spending eventually leads to added business activity and even more spending.
(b) C = 100 + 0.6 x (Y – T), I = 50, G = T = 200, NX = 0.
(i) The equilibrium level of GDP is:
Y = C + I + G = 100 + 0.6(Y – 200) + 50 + 200,
Y = 0.6Y - 120 + 350,
0.4Y = 230,
Y = 575.
(ii) The equilibrium level of GDP when the government decreases its expenditure from 200 to 150 is:
Y = C + I + G = 100 + 0.6(Y – 200) + 50 + 150,
Y = 0.6Y - 120 + 300,
0.4Y = 180,
Y = 450.
(iii) If the government attempts to achieve an output (GDP) of 1,000, then the required level of government expenditure is:
1,000 = 100 + 0.6(1000 – 200) + 50 + G,
G = 1,000 - 630 = 370.